Which of these images represents the following system of equations being solved graphically?
6x – 6y = -48
-10x + 4y = 92

Answer by mananth(16949)

6 x -6 y = -48 .............1
-10 x + 4 y = 92 .............2

equation1
when x= 0 y= 8 ( 0 , 8 )
when y = 0 x= -8 ( -8 , 0 )
Equation 2
when x= 0 y= 23 ( 0 , 23 )
when y = 0 x= -9.2 ( -9.2 , 0 )